OpenStudy (rose4825):

14x^2 – 8 + 5x – 6x^2 + 2x Which expresses the polynomial in standard form? A. 8x^2 + 7x – 8 B. 8x^2 + 3x – 8 C. 20x^2 + 7x – 8 D. 15x^2 – 8 My answer: A. 8x^2 + 7x – 8

OpenStudy (rose4825):

@sshayer

OpenStudy (eliesaab):

14x^2- 6x^2 =? 5x + 2x=? – 8 =?

OpenStudy (eliesaab):

Are you with me?

OpenStudy (eliesaab):

14x^2- 6x^2 =8x^2 5x + 2x=7x – 8 =-8

OpenStudy (eliesaab):

Your answer A is correct

OpenStudy (rose4825):

Sorry, I was still doing it on my own paper, but yes I can see how you did it. I have done all the questions already, I just need someone to confirm if I am wrong or right. :)

OpenStudy (rose4825):

Thank you!

OpenStudy (eliesaab):

YW